What is the formula for calculating the number the spectral lines in an hydrogen atom.
and also how to solve this question,
How many spectral lines are seen for hydrogen atom when electron jups from n2 =5 and n1 =1 in visible region?
 

Grade:12

6 Answers

Khimraj
7 years ago
If electron jump from nth orbital to 1st orbital then formula is n(n-1)/2. And your answer is 10.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer
megha
7 years ago
but here they have asked in the visibe region …..........
so should we consider only the balmer series ie n=2?
 
Khimraj
7 years ago
Yes absolutely for visible region you should take n =2.Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer
Debjit Mistry
6 years ago
Possible spectral lines= (n2-n1)(n2-n1+1)/2 this is the possible numbers.But in visible region, electron when jumps from n2=5 to n1=1 then according to the formula possible lines=(5-1)(5-1+1)/2=10So the possibilities are 5 to 4, 5 to 3, 5 to 2, 5 to 1, 4 to 3, 4 to 2, 4 to 1, 3 to 2, 3 to 1, 2 to 1. But here visible regin is said,i.e balmer series. so x=2 & y=3,4,5.So the possibilities are 5 to 2, 4 to 2, 3 to 2. i.e. 3 spectral lines.
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached answer to the your question.
 
The formula for finding the number of spectral lines, when an electron jumps from n2 orbit to n1 orbit is
(n2-n1)(n2-n1+1)/2
For visible spectrum n1 = 2
Hence only 3 spectral lines are possible,viz., 5 → 2, 4 → 2, 3 → 2
 
Hope it helps.
Thanks and regards,
Kushagra
Ishaan Singhal
3 years ago
orbit isn1 orbit to n2The formula for finding the number of spectral lines, when an electron jumps from
(n2-n1)(n2-n1+1)/2
 
you asked number of visible spectral lines 
therefore n2 = 5 and n1 = 2
according to the formula ,
[(5-2)(5-2+1)]/2
{3*4}/2 = 6
so therefore , 6 number of visible spectral lines will be seen for the hydrogen atom when electron jumps from n2 orbit to n1 orbit .
 
HOPE IT HELPS 
REGARDS – ISHAAN SINGHAL
PLEASE APPROVE THE ANSWER